points (1,4) and (4,2) lie on line k. what is the slope of the line that is perpendicular to k?
Answer by lwsshak3(11628)

points (1,4) and (4,2) lie on line k. what is the slope of the line that is perpendicular to k?
slope of straight line=∆y/∆x
slope of line k=(2-4)/(4-1)=-2/3
slope of line perpendicular to line k
=negative reciprocal of slope of line k
=3/2